Project

General

Profile

05 12 14

Attending: Nate, Dominick, Chris, Craig, Michael, Eric, Kanika, Satish, Adam,

Priorities: MC

— Tag with new root/art for improved memory handling?
—> Move as soon as people are ready — maybe next week?
— NuTools? 1.1.7 – will move to the new version with the new version of Art. Talk to Brian…

— Files waiting on tape — huge backlog ENSTORE can only write ~28,000 per day ( we were creating 100,000)

— Request from Andrew:
—> number of files, file sizes, and throughput requirements.
—> for different types of files.

Need to update this page:
https://cdcvs.fnal.gov/redmine/projects/nova_sam/wiki/Official_Datasets

Datasets provided for neutrinos:

Slides from Dominik:
— 11273 — Production status

— 11272: FTS ideas
— FTS proposal
— feature request – named strings
— prioritizing datasets has been important

Data:
———

- Calib:
- PCHits for calibration almost caught up
- NuMI:
- All Raw2Root caught up
- Cosmic:
- 8x10^7 ( ~4x originial request of 2x10^7 triggers).
- Reco, PID, LEM, CAF ?

Monte Carlo:
——————

— MC FD Cosmic 14 DB:
— Only 1/2 of the planned sample due to long generation time.
— 1000 s — 2x10^6 windows (500 us)
— Reco, PID, LEM, CAF (?)
— MC FD GENIE:
— 8.3x10^6 triggers (~ 300 x 3 years NOvA ?)
— Reco, PID, LEM, CAF ?

Recent request 4DB FD Cosmic MC:
——————————
— MC FD Cosmic 4DB:
— Gen, Reco, PID , LEM today…
— Run on and offsite!
— All Reco finished over the weekend.
— PID started… Nate submitted again… — status 3000 cosmic PID files finished.
— Nate and Eric submitted some jobs, —
— PID part files in SAM — LEMSUM files are not in SAM.
— Getting LEM files declared should be a priority.
—> (Michael/Robert ???)

Some production statistics:
-----------------------------------
~ 100 TB of files produced. (Estimated to be close to 200 TB with ND beam)
~ 100,000 files per day being sent to tape (tape can only do 28,000)

20k CPU days (dominated by FD cosmics PID?)

Average 800 jobs running for the last month on site (MC gen offsite)